Accountancy

The average profit of a firmfor the last five years is rs 20000 and the capital invested is rs 130000.the normal profit of the similar line of a business is 10% . Calculate the amount of goodwill if it is 2 years purchase of super profit 
Answer #1

Average profit = 20000

Normal profit=130000×10/100=13000

Super profit = 20000-13000=7000

Goodwill=7000×2=14000

answered by: Advika